### v2

# Chow and augmented Chow polynomials as evaluations of Poincaré-extended $\mathbf{a}\mathbf{b}$-indices

Abstract.
We show that Chow polynomials and augmented Chow polynomials of matroids, and more generally of finite graded posets admitting $R$-labellings, are obtained as evaluations of their Poincaré-extended $\mathbf{a}\mathbf{b}$-indices. This implies in particular explicit combinatorial $\gamma$-positive expansions for both. We then evaluate this expansion to obtain a compact expansion for the braid arrangement.
2020 Mathematics Subject Classification:
Primary 06A07; Secondary 05B35, 52C40
1. Main results
The Chow ring $\underline{\textsf{CH}}(M)$ and the augmented Chow ring $\textsf{CH}(M)$ are important geometric-algebraic object attached to a loopless matroid $M$ that have been introduced in [FY04] and in [BHM+22], respectively. They have attracted a lot of attention in recent years, maybe most notable in their celebrated use in the proof of the log-concavity of the Whitney numbers of the second kind in [AHK18]. We follow [FMSV24] in our notations and call their Hilbert-Poincaré series $\underline{\textsf{H}}_{M}(x)$ and $\textsf{H}_{M}(x)$ the Chow polynomial and the augmented Chow polynomial.
To state the two main results, recall that the lattice of flats $\mathcal{L}(M)$ of the matroid $M$ is geometric and thus admits an $R$-labeling $\lambda$, i.e., an edge-labeling $\lambda$ such that each interval in $\mathcal{L}(M)$ contains a unique maximal chain along which $\lambda$ is increasing [Sta74]. To a maximal chain =$\mathcal{F}=\{\mathcal{F}_{0}\prec\dots\prec\mathcal{F}_{n}\}$ in $\mathcal{L}(M)$ associate the sequence $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ of edge labels $\lambda_{i}=\lambda(\mathcal{F}_{i-1}\prec\mathcal{F}_{i})$ for $n={\sf rank}(M)$ being the rank of $M$.
Theorem 1.1.
The Chow polynomial $\underline{\textsf{H}}_{M}(x)$ has the expansion
$\underline{\textsf{H}}_{M}(x)$ $=\sum_{\mathcal{F}}x^{\mathsf{des}(\lambda_{\mathcal{F}})}(x+1)^{n-1-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,,$
where the sum ranges over all maximal chains $\mathcal{F}$ in $\mathcal{L}(M)$ such that for $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ it holds that $\lambda_{1}<\lambda_{2}$ and also that $\lambda_{i}>\lambda_{i+1}$ implies $\lambda_{i-1}<\lambda_{i}$ for $i\in\{2,\dots,n-1\}$.
Theorem 1.2.
The augmented Chow polynomial $\textsf{H}_{M}(x)$ has the expansion
$\textsf{H}_{M}(x)$ $=\sum_{\mathcal{F}}x^{\mathsf{des}(\lambda_{\mathcal{F}})}(x+1)^{n-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,,$
where the sum ranges over all maximal chains $\mathcal{F}$ in $\mathcal{L}(M)$ such that for $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ it holds that $\lambda_{i}>\lambda_{i+1}$ implies $\lambda_{i-1}<\lambda_{i}$ for $i\in\{2,\dots,n-1\}$.
Corollary 1.3.
The Chow polynomial $\underline{\textsf{H}}_{M}(x)$ and the augmented Chow polynomial $\textsf{H}_{M}(x)$ are $\gamma$-positive. They are in particular nonnegative, palindromic and unimodal.
To the best of our knowledge, the only known proof of the $\gamma$-positivity appears in [FMSV24] and relies on semi-small decompositions for the Chow and the augmented Chow rings [BHM+22] or, equivalently, on the Kähler package for these rings [AHK18]. Hence, Theorems 1.1, LABEL: and 1.2 may be viewed as a Hodge-theory free proof to the $\gamma$-positivity and also of unimodality of Chow and augmented Chow polynomials of geometric lattices. After proving the main theorems in Section 2, we evaluate the expansions in the case of the braid arrangement in Section 3.
Remark 1.4 (Generalization to finite graded posets admitting $R$-labelings).
In forthcoming work by Ferroni–Matherne–Vecchi, the definitions of the Chow and augmented Chow polynomials are generalized to finite graded posets with $\hat{0}$ and $\hat{1}$ [FMV24], see Definition 2.1. Using this definition for general finite graded posets, our proofs hold for all such posets admitting $R$-labelings. We in particular obtain non-negativity, palindromicity, unimodality, and $\gamma$-positivity in this generality.
To prove the main results, we show for finite graded posets admitting $R$-labelings that the Chow and the augmented Chow polynomials are obtained as evaluations of the Poincaré-extended $\mathbf{a}\mathbf{b}$-index defined and studied by the author and Dorpalen-Barry–Maglione in [DMS23]. It turns out that the Chow polynomial, but not the augmented Chow polynomial, is an evaluation of the coarse flag Hilbert-Poincaré series as defined and studied by Maglione–Voll in [MV23], which itself is also an evaluation of the extended $\mathbf{a}\mathbf{b}$-index. As the Poincaré-extended $\mathbf{a}\mathbf{b}$-indices of finite graded poset admitting $R$-labelings have explicit descriptions using the labeling along maximal chains, we then deduce Theorems 1.1, LABEL: and 1.2.
Remark 1.5 (Three existing real-rootedness conjectures).
We show in Theorem 2.6 that the Chow polynomial is an evaluation of the extended $\mathbf{a}\mathbf{b}$-index. This evaluation is conjectured to be real-rooted [FS23, Conjecture 8.18]. It is shown in [DMS23, Corollary 2.23] that the numerator polynomial of the coarse flag Hilbert-Poincaré series is an evaluation of the extended $\mathbf{a}\mathbf{b}$-index and asked in [KM23, Question 1.6] for which matroids a further evaluation is real-rooted. Also the chain polynomial of the lattice of flats is yet another evaluation of the extended $\mathbf{a}\mathbf{b}$-index which is conjectured in [AKE23, Conjecture 1.2] to be real-rooted. This suggests the following question.
Question 1.6.
Which evaluations of the extended $\mathbf{a}\mathbf{b}$-index of a matroid, or of a more general finite graded poset admitting an $R$-labeling, are real-rooted?
Acknowledgment
This work was conducted at the Mathematisches Forschungsinstitut Oberwolfach in June 2024 during the workshop “Arrangements, Matroids and Logarithmic Vector Fields”. We very much thank the institute for their hospitality and the organizers Takuro Abe, Graham Denham, Eva-Maria Feichtner, and Gerhard Röhrle for setting up this wonderful and inspiring meeting. The results in this paper were obtained after attending the talk by Luis Ferroni entitled “Chow polynomials of posets”. The author would very much like to thank Luis and his collaborators Chris Eur, Jacob Matherne, and Lorenzo Vecchi, and as well Elena Hoster and Lukas Kühne for inspiring discussions concerning Chow polynomials and their combinatorial interpretations, and for comments on a preliminary version of this paper.
2. The Poincaré-extended $\mathbf{a}\mathbf{b}$-index and its evaluations
Let $P$ be a finite graded poset of rank $n$. That is, $P$ is a finite poset with unique minimum element $\hat{0}$ of rank 00 and unique maximum element $\hat{1}$ of rank $n$ such that ${\sf rank}(F)$ is equal to the length of any maximal chain from $\hat{0}$ to $F$. Its Möbius function $\mu$ is given by $\mu(F,F)=1$ for all $F\in P$ and $\mu(F,G)=-\sum_{F\leq H<G}\mu(F,H)$ for all $F<G$. Its Poincaré polynomial is
$\mathcal{\sf Poin}_{P}(q)$ $=\sum_{F\in P}\mu(\hat{0},F)\cdot(-q)^{{\sf rank}(F)}\,,$
its characteristic polynomial is
$\chi_{P}(q)$ $=\sum_{F\in P}\mu(\hat{0},F)\cdot q^{n-{\sf rank}(F)}\,,$
and its reduced characteristic polynomial is
$\overline{\chi}_{P}(q)$ $=\chi_{P}(q)/(q-1)=q^{n}\cdot\mathcal{\sf Poin}_{P}(-q^{-1})/(q-1)\,.$
A (not necessary maximal) chain $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k}<\mathcal{C}_{k+1}\}$ in $P$ is an ordered set of pairwise comparable elements. We assume throughout this paper that all chains end in the maximum element $\mathcal{C}_{k+1}=\hat{1}$. A chain is maximal if $k=n$ and we usually write $\{\mathcal{F}_{0}\prec\dots\prec\mathcal{F}_{n}\}$ in this case in order to clearly distinguish between general chains ending in $\hat{1}$ and maximal chains. The chain Poincaré polynomial and the reduced chain characteristic polynomial of such a chain $\mathcal{C}$ are
$\mathcal{\sf Poin}_{P,\mathcal{C}}(q)=\prod_{i=1}^{k}\mathcal{\sf Poin}_{[\mathcal{C}_{i},\mathcal{C}_{i+1}]}(q)\quad\text{ and }\quad\overline{\chi}_{P,\mathcal{C}}(q)=\prod_{i=1}^{k}\overline{\chi}_{[\mathcal{C}_{i},\mathcal{C}_{i+1}]}(q)\,,$
where $[\mathcal{C}_{i},\mathcal{C}_{i+1}]\subseteq P$ denotes the interval between two consecutive elements in the chain. In particular,
$\mathcal{\sf Poin}_{P,\{\hat{0}<\hat{1}\}}(q)=\mathcal{\sf Poin}_{P}(q),\quad\overline{\chi}_{P,\{\hat{0}<\hat{1}\}}(q)=\overline{\chi}_{P}(q)\,.$
2.1. The Chow and the augmented Chow polynomials
The following description of the Chow and augmented Chow polynomials of a matroid is taken from [FMSV24, Theorem 1.3 & 1.4]. We present it in the generality of general finite graded posets as introduced by [FMV24].
Definition 2.1 (Ferroni–Matherne–Vecchi).
Let $P$ be a finite graded poset. The Chow polynomial is defined to be $1$ if ${\sf rank}(P)=0$ and otherwise
$\underline{\textsf{H}}_{P}(x)=\sum_{\mathcal{C}}\overline{\chi}_{P,\mathcal{C}}(x)\ \in\mathbb{Z}[x]\,,$
where the sum ranges over all chains $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k+1}\}$ in $P$ starting at $\mathcal{C}_{1}=\hat{0}$ and ending in $\mathcal{C}_{k+1}=\hat{1}$. The augmented Chow polynomial is defined to be $1$ if ${\sf rank}(P)=0$ and otherwise
$\textsf{H}_{P}(x)=\sum_{\mathcal{C}}x^{{\sf rank}(\mathcal{C}_{1})}\cdot\overline{\chi}_{P,\mathcal{C}}(x)\ \in\mathbb{Z}[x]\,,$
where the sum ranges over all chains $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k+1}\}$ in $P$ ending in $\mathcal{C}_{k+1}=\hat{1}$.
It is shown in [FMSV24, Theorems 1.3 & 1.4] that for the lattice of flats $P=\mathcal{L}(M)$, this definition agrees with the Hilbert-Poincaré series of the Chow and the augmented Chow rings for a loopless matroid $M$ as presented in Section 1.
2.2. The Poincaré-extended $\mathbf{a}\mathbf{b}$-index
The following description and properties of the extended $\mathbf{a}\mathbf{b}$-index are taken from [DMS23]. Throughout this section, let $P$ be a finite graded poset admitting an $R$-labeling, i.e., an edge-labeling $\lambda:\mathcal{E}(P)\rightarrow\mathbb{N}_{+}$ such that each interval in $P$ contains a unique maximal chain along which $\lambda$ is (weakly) increasing.
Let $\mathbb{Z}[y]\langle\mathbf{a},\mathbf{b}\rangle$ be the polynomial ring in two noncommuting variables $\mathbf{a},\mathbf{b}$ with coefficients being polynomials in the variable $y$. For a subset $S\subseteq\{0,\dots,n-1\}$, we set ${\sf wt}_{S}(\mathbf{a},\mathbf{b})=w_{0}\cdots w_{n-1}$ for the polynomial with
(1) $w_{k}=\begin{cases}\mathbf{b}&\text{if }k\in S,\\ \mathbf{a}-\mathbf{b}&\text{if }k\notin S\,.\end{cases}$
For a chain $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k}<\mathcal{C}_{k+1}\}$ in $P$, ending in $\mathcal{C}_{k+1}=\hat{1}$, we moreover set
${\sf wt}_{\mathcal{C}}={\sf wt}_{\{{\sf rank}(\mathcal{C}_{1}),\dots,{\sf rank}(\mathcal{C}_{k})\}}\,.$
Definition 2.2.
The (Poincaré-)extended $\mathbf{a}\mathbf{b}$-index of $P$ is
${\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})=\sum_{\mathcal{C}}\mathcal{\sf Poin}_{P,\mathcal{C}}(y)\cdot{\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})\ \in\mathbb{Z}[y]\langle\mathbf{a},\mathbf{b}\rangle\,,$
where the sum ranges over all chains $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k+1}\}$ in $P$ ending in $\mathcal{C}_{k+1}=\hat{1}$.
Since $P$ has a unique minimum, we always have $\mathcal{\sf Poin}_{P}(0)=1$, implying
(2) ${\text{\scriptsize\sf ex}}\Psi_{P}(0,\mathbf{a},\mathbf{b})=\Psi_{P}(\mathbf{a},\mathbf{b})=\sum_{\mathcal{C}}{\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})$
for (a mild variation of) the $\mathbf{a}\mathbf{b}$-index $\Psi_{P}(\mathbf{a},\mathbf{b})$ as given, for example, in [Bay21, Section 2].
Proposition 2.3 ([DMS23, Corollary 2.22]).
We have
$\mathbf{b}\cdot\iota\big{(}{\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})\big{)}=\sum_{\mathcal{C}}\mathcal{\sf Poin}_{P,\mathcal{C}}(y)\cdot{\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})\ \in\mathbb{N}[y]\langle\mathbf{a},\mathbf{b}\rangle\,,$
where the sum ranges over all chains $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k+1}\}$ starting in $\mathcal{C}_{1}=\hat{0}$ and ending in $\mathcal{C}_{k+1}=\hat{1}$ and where $\iota$ deletes the first variable from every $\mathbf{a}\mathbf{b}$-monomial.
Observe that the factor $\mathbf{b}$ on the left-hand side comes from the assumption that $\mathcal{C}_{1}=\hat{0}$ and thus, every $\mathbf{a},\mathbf{b}$-polynomial ${\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})$ starts with the variable $\mathbf{b}$ as $0={\sf rank}(\mathcal{C}_{1})$. For later reference, we set
(3) ${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})=\iota\big{(}{\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})\big{)}\,.$
We next present two of the main result of [DMS23], from which we then deduce the proposed evaluation to obtain the Chow and the augmented Chow polynomials. As in Section 1, we write $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ for the labeling along a maximal chain =$\mathcal{F}=\{\mathcal{F}_{0}\prec\dots\prec\mathcal{F}_{n}\}$ with $\lambda_{i}=\lambda(\mathcal{F}_{i-1}\prec\mathcal{F}_{i})$. For a subset $E\subseteq\{1,\dots,n\}$, we moreover set $\lambda_{\mathcal{F},E}=(\lambda^{\prime}_{0},\lambda^{\prime}_{1},\dots,\lambda^{\prime}_{n})$ with $\lambda^{\prime}_{0}=0$ and $\lambda^{\prime}_{i}=-\lambda_{i}$ if $i\in E$ and $\lambda^{\prime}_{i}=\lambda_{i}$ if $i\notin E$. We in addition define ${\sf u}_{\mathcal{F},E}(\mathbf{a},\mathbf{b})=u_{0}\cdots u_{n-1}$ to be the monomial in $\mathbf{a},\mathbf{b}$ with
$u_{i}=\begin{cases}\mathbf{a}&\text{if }\lambda^{\prime}_{i}\leq\lambda^{\prime}_{i+1}\,,\\ \mathbf{b}&\text{if }\lambda^{\prime}_{i}>\lambda^{\prime}_{i+1}\,,\\ \end{cases}$
so the monomial ${\sf u}_{\mathcal{F},E}(\mathbf{a},\mathbf{b})$ detects the ascent-descent pattern of $\lambda_{\mathcal{F},E}$.
Theorem 2.4 ([DMS23, Theorem 2.7]).
Let $P$ be a finite graded poset of rank $n$, admitting an $R$-labeling $\lambda$. Then
${\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})=\sum_{\mathcal{F},E}y^{\#E}\cdot{\sf u}_{\mathcal{F},E}(\mathbf{a},\mathbf{b})$
where the sum ranges over all maximal chains $\mathcal{F}$ in $P$ and all subsets $E\subseteq\{1,\dots,n\}$.
Corollary 2.5 ([DMS23, Corollary 2.9]).
Let $P$ be a finite graded poset of rank $n$, admitting an $R$-labeling $\lambda$. Then
${\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})=\omega\big{(}\Psi_{P}(\mathbf{a},\mathbf{b})\big{)}$
where the substitution $\omega$ first replaces all occurrences of $\mathbf{a}\mathbf{b}$ with $\mathbf{a}\mathbf{b}+y\mathbf{b}\mathbf{a}+y\mathbf{a}\mathbf{b}+y^{2}\mathbf{b}\mathbf{a}$ and then simultaneously replaces all remaining occurrences of $\mathbf{a}$ with $\mathbf{a}+y\mathbf{b}$ and $\mathbf{b}$ with $\mathbf{b}+y\mathbf{a}$.
It is immediate that the $\mathbf{a}\mathbf{b}$-index in (2) has the expansion
(4) $\Psi_{P}(\mathbf{a},\mathbf{b})=\sum_{\mathcal{F}}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})$
and that every monomial ${\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})=u_{0}\cdots u_{n-1}$ starts with the variable $u_{0}=\mathbf{a}$.
2.3. Evaluating the Poincaré-extended $\mathbf{a}\mathbf{b}$-index
The main results of this section are the following evaluations.
Theorem 2.6.
Let $P$ be a finite graded poset of rank $n$ admitting an $R$-labeling. Then
${\text{\scriptsize\sf ex}}\Psi_{P}(-x,1,x)$ $=(1-x)^{n}\cdot\textsf{H}_{P}(x)\,,$
${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,1,x)$ $=(1-x)^{n}\cdot\underline{\textsf{H}}_{P}(x)\,.$
Before proving this evaluation, we deduce Theorems 1.1, LABEL: and 1.2.
Proof of Theorems 1.1, LABEL: and 1.2.
Using the expansion (4), an immediate consequence of Theorem 2.4 and Corollary 2.5 is that
${\text{\scriptsize\sf ex}}\Psi_{P}(-x,1,x)=\sum_{\mathcal{F}}\omega\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{|}_{y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x}$
where the sum ranges over all maximal chains $\mathcal{F}$ in $P$. Applying the evaluation $y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x$ to the substitution $\omega$ gives that every occurrence in ${\sf u}_{\mathcal{F},\emptyset}$ of the pattern $\mathbf{a}\mathbf{b}$ becomes a factor $x-x^{2}-x^{2}+x^{3}=x(1-x)^{2}$, every remaining $\mathbf{a}$ yields a factor $1-x^{2}=(1+x)(1-x)$, and every remaining $\mathbf{b}$ yields a factor 00. We therefore obtain
$\omega\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{|}_{y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x}$ $=0$
if $\lambda_{\mathcal{F}}$ contains a descent that does not immediately follow an ascent, and otherwise
$\omega\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{|}_{y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x}$ $=(1-x)^{n}\cdot x^{\mathsf{des}(\lambda_{\mathcal{F}})}(1+x)^{n-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,.$
Summing over all maximal chains $\mathcal{F}$ in $P$ such that for $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ it holds that $\lambda_{i}>\lambda_{i+1}$ implies $\lambda_{i-1}<\lambda_{i}$ for $i\in\{2,\dots,n-1\}$ then gives
${\text{\scriptsize\sf ex}}\Psi_{P}(-x,1,x)=(1-x)^{n}\sum_{\mathcal{F}}x^{\mathsf{des}(\lambda_{\mathcal{F}})}(x+1)^{n-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,.$
This proves Theorem 1.2.
The argument to deduce Theorem 1.1 is almost the same. The only difference is that now every initial variable is removed from every $\mathbf{a}\mathbf{b}$-monomial to obtain ${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})$ from ${\text{\scriptsize\sf ex}}\Psi_{P}(x,\mathbf{a},\mathbf{b})$. This is, we have
${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})=\iota\big{(}{\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})\big{)}=\sum_{\mathcal{F}}\iota\big{(}\omega\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{)}\,.$
We consider the two cases whether ${\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})$ starts with $\mathbf{a}\mathbf{a}$ or with $\mathbf{a}\mathbf{b}$. If it starts with $\mathbf{a}\mathbf{a}$, we obtain
$\iota\big{(}\omega\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{)}=(1+y)\omega\big{(}\iota\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{)}$
as the first variable $\mathbf{a}$ is evaluated to $\iota\omega(\mathbf{a})=\iota(\mathbf{a}+y\mathbf{b})=1+y$. If it starts with $\mathbf{a}\mathbf{b}$, we obtain
$\iota\omega\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{|}_{y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x}=0$
as the initial two variables $\mathbf{a}\mathbf{b}$ are evaluated to
$\iota\omega(\mathbf{a}\mathbf{b})=\iota((1+y)\mathbf{a}\mathbf{b}+y(1+y)\mathbf{b}\mathbf{a})=(1+y)\mathbf{b}+y(1+y)\mathbf{a}$
and
$(1+y)\mathbf{b}+y(1+y)\mathbf{a}\big{|}_{y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x}=0$
Summing now over all maximal chains $\mathcal{F}$ in $P$ such that for $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ it holds that $\lambda_{1}<\lambda_{2}$ and also that $\lambda_{i}>\lambda_{i+1}$ implies $\lambda_{i-1}<\lambda_{i}$ for $i\in\{2,\dots,n-1\}$ then gives
${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,1,x)=(1-x)^{n}\sum_{\mathcal{F}}x^{\mathsf{des}(\lambda_{\mathcal{F}})}(x+1)^{n-1-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,.$
This proves Theorem 1.1. ∎
Proof of Theorem 2.6.
We evaluate the Chow polynomial and the extended $\mathbf{a}\mathbf{b}$-index as follows, where the sums range over all chains $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k}<\mathcal{C}_{k+1}\}$ starting in $\mathcal{C}_{1}=\hat{0}$ and ending in $\mathcal{C}_{k+1}=\hat{1}$. We write $k_{\mathcal{C}}=k$ to denote its length. Then
$\underline{\textsf{H}}_{P}(x)$ $=\sum_{\mathcal{C}}\overline{\chi}_{P,\mathcal{C}}(x)=\sum_{\mathcal{C}}\prod_{i=1}^{k_{\mathcal{C}}}\overline{\chi}_{[\mathcal{C}_{i},\mathcal{C}_{i+1}]}(x)$
$=\sum_{\mathcal{C}}\prod_{i=1}^{k_{\mathcal{C}}}x^{{\sf rank}(\mathcal{C}_{i+1})-{\sf rank}(\mathcal{C}_{i})}(x-1)^{-1}\cdot\mathcal{\sf Poin}_{[\mathcal{C}_{i},\mathcal{C}_{i+1}]}(-x^{-1})$
$=x^{n}\sum_{\mathcal{C}}\prod_{i=1}^{k_{\mathcal{C}}}(x-1)^{-1}\cdot\mathcal{\sf Poin}_{[\mathcal{C}_{i},\mathcal{C}_{i+1}]}(-x^{-1})$
$=x^{n}\sum_{\mathcal{C}}(x-1)^{-k_{\mathcal{C}}}\mathcal{\sf Poin}_{P,\mathcal{C}}(-x^{-1})$
$=x^{n}(x-1)^{-1}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x^{-1},\tfrac{x}{x-1},\tfrac{1}{x-1})$
where we used in the last equation that ${\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})$ has exactly $n-1$ factors, either being $\mathbf{b}$ for ranks in $\mathcal{C}\setminus\{\mathcal{C}_{k+1}\}$ or being $\mathbf{a}-\mathbf{b}$ for ranks not in $\mathcal{C}$, and moreover that evaluating $\mathbf{b}\mapsto\frac{1}{x-1}$ and $\mathbf{a}\mapsto\frac{x}{x-1}$ evaluates $\mathbf{a}-\mathbf{b}\mapsto 1$ and at the same time yields a factor $(x-1)^{1-k_{\mathcal{C}}}$. Using the symmetry $\underline{\textsf{H}}_{P}(x)=x^{n-1}\underline{\textsf{H}}_{P}(x^{-1})$, we obtain
$\underline{\textsf{H}}_{P}(x)$ $=x^{-1}(x^{-1}-1)^{-1}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,\tfrac{1}{1-x},\tfrac{x}{1-x})$
$=(1-x)^{-n}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,1,x)$
where we used that ${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})$ is homogeneous of $\mathbf{a}\mathbf{b}$-degree $n-1$ and thus
${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,\tfrac{1}{1-x},\tfrac{x}{1-x})=(1-x)^{n-1}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,1,x)\,.$
The argument for
$\textsf{H}_{P}(x)=(1-x)^{-n}{\text{\scriptsize\sf ex}}\Psi_{P}(-x,1,x)$
is word for word the same with the only difference that chains $\mathcal{C}=\{\mathcal{C}_{1}<\dots<\mathcal{C}_{k}<\mathcal{C}_{k+1}\}$ do not necessarily need to start in $\mathcal{C}_{1}=\hat{0}$, and thus the factor $x^{{\sf rank}(\mathcal{C}_{1})}$ in the definition of the augmented Chow polynomial ensures that $x^{n}=x^{{\sf rank}(\mathcal{C}_{1})}\cdot\prod_{i=1}^{k_{\mathcal{C}}}x^{{\sf rank}(\mathcal{C}_{i+1})-{\sf rank}(\mathcal{C}_{i})}$. ∎
Remark 2.7.
We used in the proof the symmetry $\underline{\textsf{H}}_{P}(x)=x^{n-1}\underline{\textsf{H}}_{P}(x^{-1})$. As the symmetry of the evaluated extended $\mathbf{a}\mathbf{b}$-index is clear from its combinatorial description, this is not a prerequired property of the Chow polynomial and we only used it here for the clarity of the argument.
2.4. Evaluating the coarse flag Hilbert-Poincaré series
The coarse flag Hilbert-Poincaré series ${\sf cfHP}_{P}(y,t)$ is originally defined and studied in [MV23] as an algebraic invariant in geometric group theory. It is shown in [DMS23, Corollary 2.22] that this rational function can be expressed as
${\sf cfHP}_{P}(y,t)=(1-t)^{-n}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,1,t)\,.$
We thus immediately obtain the following specialization from Theorem 2.6.
Corollary 2.8.
Let $P$ be a finite graded poset of rank $n$ admitting an $R$-labeling. The Chow polynomial is then obtained from the coarse flag Hilbert-Poincaré series by the specialization
$\underline{\textsf{H}}_{P}(x)={\sf cfHP}_{P}(x,-x)\,.$
Applying $\leavevmode\nobreak\ \iota$ to ${\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})$ restricts the situation to chains starting in $\hat{0}$ as in Proposition 2.3. It is thus not immediately possible to also obtain the augmented Chow polynomial from the coarse flag Hilbert-Poincaré series.
3. Chow and augmented Chow polynomials for braid arrangement
In this section, we apply the main theorems to obtain a compact expansion of the Chow and the augmented Chow polynomials for the braid arrangement $\mathcal{B}_{n}$ given by the hyperplanes
[…]
in $V=\big{\{}(x_{1},\dots,x_{n+1})\in\mathbb{R}^{n+1}\mid x_{1}+\dots+x_{n+1}=0\big{\}}\cong\mathbb{R}^{n}$.
Theorem 3.1.
We have
$\underline{\textsf{H}}_{\mathcal{B}_{n}}(x)$ $=\sum_{(a_{1},\dots,a_{n})}a_{1}\cdot\ldots\cdot a_{n}\cdot x^{\mathsf{des}(a_{1},\dots,a_{n})}(x+1)^{n-1-2\ \!\!\mathsf{des}(a_{1},\dots,a_{n})}\,,$
where the sum ranges over all tuples $(a_{1},\dots,a_{n})$ with $a_{i}\in\{1,\dots,n+1-i\}$ such that $a_{1}\leq a_{2}$ and $a_{i}>a_{i+1}$ implies $a_{i-1}\leq a_{i}$ for $i\in\{2,\dots,n-1\}$. We moreover have
$\textsf{H}_{\mathcal{B}_{n}}(x)$ $=\sum_{(a_{1},\dots,a_{n})}a_{1}\cdot\ldots\cdot a_{n}\cdot x^{\mathsf{des}(a_{1},\dots,a_{n})}(x+1)^{n-2\ \!\!\mathsf{des}(a_{1},\dots,a_{n})}\,,$
where the sum ranges over all tuples $(a_{1},\dots,a_{n})$ with $a_{i}\in\{1,\dots,n+1-i\}$ such that $a_{i}>a_{i+1}$ implies $a_{i-1}\leq a_{i}$ for $i\in\{2,\dots,n-1\}$.
Example 3.2.
Let $n=2$. For the Chow polynomials, the sum ranges over $\{(1,1)\}$ and we obtain
$\underline{\textsf{H}}_{\mathcal{B}_{2}}(x)$ $=1\cdot 1\cdot x^{0}(x+1)^{2-1-0}=x+1$
and for the augmented Chow polynomials, the sum ranges over $\{(1,1),(2,1)\}$ and we obtain
$\textsf{H}_{\mathcal{B}_{2}}(x)$ $=1\cdot 1\cdot(x+1)^{2}+1\cdot 2\cdot x^{1}=x^{2}+4x+1\,.$
Let now $n=3$. For the Chow polynomials, the sum ranges over $\{(1,1,1),(1,2,1),(2,2,1)\}$ and we obtain
$\underline{\textsf{H}}_{\mathcal{B}_{3}}(x)$ $=1\cdot(x+1)^{2}+2\cdot x+4\cdot x=x^{2}+8x+1$
[…]
$\textsf{H}_{\mathcal{B}_{3}}(x)$ $=1\cdot(x+1)^{3}+2\cdot x(x+1)+4\cdot x(x+1)+2\cdot x(x+1)+3\cdot x(x+1)$
$=(x+1)^{3}+11\cdot x(x+1)$
$=x^{3}+14x^{2}+14x+1\,.$
We prove Theorem 3.1 by analyzing the maximal chains in the lattice of flats of the braid arrangement for one particular $R$-labeling. The lattice of flats of $\mathcal{B}_{n}$ is the lattice of set partitions of $\{1,\dots,n+1\}$ of rank $n$, with cover relations being given by merging two blocks, denoted by $\Pi_{n}$. Its minimal element is the all-singletons set partition and its maximal element is the all-in-one-block set partition. It moreover has an $R$-labeling $\lambda$ given by labeling a cover relation in which the two blocks $B$ and $B^{\prime}$ are merged by the max-of-min labeling $\max\big{\{}\!\min(B),\min(B^{\prime})\big{\}}$, see [Bjö80, Example 2.9]. Observe that the max-of-min labeling along any maximal chain is a permutation of $\{2,\dots,n+1\}$.
Remark 3.3.
This $R$-labeling was recently used in [HS24] to derive an explicit form for the coarse flag Hilbert-Poincaré series of the braid arrangement. In light of Corollary 2.8, it seems natural that the max-to-min labeling also plays a crucial role in the understanding of the Chow polynomials of the braid arrangements.
To deduce Theorem 3.1 from Theorems 1.1, LABEL: and 1.2, we calculate for a given permutation $\sigma$ of $\{2,\dots,n+1\}$ the number of maximal chains $\mathcal{L}$ in $\Pi_{n}$ for which $\lambda_{\mathcal{F}}=\sigma$. (Our taylormade version of) the inversion sequence $\mathsf{Inv}(\sigma)=(a_{1},\dots,a_{n})$ of a permutation $\sigma=[\sigma_{1},\dots,\sigma_{n}]$ of $\{2,\dots,n+1\}$ in one-line notation is given by
$a_{i}=\#\{j\geq i\mid\sigma_{j}\leq\sigma_{i}\}\,.$
Observe that the entry $a_{i}$ in the inversion sequence $\mathsf{Inv}(\sigma)$ may be computed from the partial permutation $[\sigma_{1},\dots,\sigma_{i}]$ as
(5) $a_{i}=\sigma_{i}-\#\{j\leq i\mid\sigma_{j}\leq\sigma_{i}\}\,.$
Example 3.4.
Let $n=3$. Then
$\sigma$ $\mathsf{Inv}(\sigma)$
$[2,3,4]$ $(1,1,1)$
$[2,4,3]$ $(1,2,1)$
$[3,2,4]$ $(2,1,1)$
$[3,4,2]$ $(2,2,1)$
$[4,2,3]$ $(3,1,1)$
$[4,3,2]$ $(3,2,1)$
The map $\sigma\mapsto\mathsf{Inv}(\sigma)$ is moreover a bijection between permutations of $\{2,\dots,n+1\}$ and sequences $(a_{1},\dots,a_{n})$ for which $a_{i}\in\{1,\dots,n+1-i\}$.
Proposition 3.5.
Let $\sigma$ be a permutation of $\{2,\dots,n+1\}$ with inversion sequence $\mathsf{Inv}(\sigma)=(a_{1},\dots,a_{n})$. Then
#$\#\{\text{maximal chains }\mathcal{F}\text{ in }\Pi_{n}\mid\lambda_{\mathcal{F}}=\sigma\}=a_{1}\cdot\ldots\cdot a_{n}\,.$
Proof.
Given a set partition $\mathbf{B}=B_{1}|\dots|B_{\ell}\in\Pi_{n}$ and a maximal chain $\mathcal{F}$ going through $\mathbf{B}$, the max-of-min labels among the cover relations in $\mathcal{F}$ before reaching $\mathbf{B}$ are those indices that have already appeared as max-of-min’s. These are
$\{2,\dots,n+1\}\setminus\{\min B_{1},\dots,\min B_{\ell}\}\,.$
Conversely, the labels after reaching $\mathbf{B}$ are $\{\min B_{1},\dots,\min B_{\ell}\}\setminus\{1\}$. Moreover, the labeled interval between $\mathbf{B}$ and the maximal element $\{1,\dots,n+1\}\in\Pi_{n}$ does not depend on $\mathbf{B}$ itself as it is simply the set partition lattice on the set $\{\min B_{1},\dots,\min B_{k}\}$ again with the max-of-min labeling.
Let now $i\in\{1,\dots,n\}$, let $\sigma=[\sigma_{1},\dots,\sigma_{i}]$ be a non-empty partial permutation, i.e., a list of $i$ distinct numbers in $\{2,\dots,n+1\}$ and let $\sigma^{\prime}=[\sigma_{1},\dots,\sigma_{i-1}]$ be the partial permutation obtained by removing from $\sigma$ the last entry. Set $a_{i}$ to be the $i$’s entry in the inversion sequence of any permutation starting with $\sigma$ as given in (5). Finally, denote by $m_{\sigma}$ the number of maximal chains $\mathcal{F}$ in $\Pi_{n}$ for which the labeling $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ starts with $(\lambda_{1},\dots,\lambda_{i})=(\sigma_{1},\dots,\sigma_{i})$ and by $m_{\sigma^{\prime}}$ the number of maximal chains for which the labeling starts with $(\sigma_{1},\dots,\sigma_{i-1})$. Then
$m_{\sigma}=m_{\sigma^{\prime}}\cdot a_{i}\,,$
because there are exactly $a_{i}$ many blocks $B$ into which one can merge the block containing $\sigma_{i}$, say $B^{\prime}$, such that $\sigma_{i}=\max\{\min B,\min B^{\prime}\}$. The statement follows from recursively applying this counting formula. ∎
Proof of Theorem 3.1.
Let $\sigma=[\sigma_{1},\dots,\sigma_{n}]$ be a permutation of $\{2,\dots,n+1\}$ with inversion sequence $\mathsf{Inv}(\sigma)=(a_{1},\dots,a_{n})$. Then $\sigma_{i}>\sigma_{i+1}$ is a descent of $\sigma$ if and only if $a_{i}>a_{i+1}$ is a descent of the inversion sequence. Both parts of the statement follow by using Proposition 3.5 to simplify the ranges of the sums in Theorems 1.1, LABEL: and 1.2. ∎